OpenStudy (anonymous):

Hint: Remember the reciprocal rule when dividing with fractions. \frac{1}{3}\div 6 A. 2 B. \frac{1}{9} C. \frac{1}{18}

OpenStudy (anonymous):

1 - divided by 6 3

OpenStudy (anonymous):

A. 2

OpenStudy (anonymous):

o.

OpenStudy (anonymous):

your welcome :^)

OpenStudy (mathstudent55):

The reciprocal rule in division is that to divide two numbers, you multiply the first number by the reciprocal of the second number. \( \large \dfrac{1}{3} \div 6\) First look at the problem and think about it. Let's say you have a third of a chocolate bar, and you want to divide it by 6 people. How much will each person get? Each person will get 1/6 of 1/3. It will be a very small piece. That means that the answer A, 2, makes no sense. How can you divide 1/3 by 6 and get 2, more than you started with? Now let's actually do it using the reciprocal rule. \( \large \dfrac{1}{3} \div 6 = \dfrac{1}{3} \times \dfrac{1}{6} = \dfrac{1 \times 1}{3 \times 6} = \dfrac{1}{18}\) This result makes sense. If you divide 1/3 of a chocolate bar by 6 people, each person gets 1/18 of the chocolate bar. james' mistake above was that he forgot the reciprocal rule and multiplied by 6 instead of multiplying by the reciprocal of 6 which is 1/6.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):